How would you choose to reduce the system shown to a 2 × 2? Explain why you would choose this approach.
–3x + y – 2z = 10 (1)
5x – 2y – 2z = 12 (2)
x – y + z = 23 (3)

Answer by Alan3354(69443)

How would you choose to reduce the system shown to a 2 × 2? Explain why you would choose this approach.
–3x + y – 2z = 10 (1)
5x – 2y – 2z = 12 (2)
x – y + z = 23 (3)
===================
I would eliminate y or z.
There's no advantage to either.
Eliminating x would require more work.